The tempreature of village reached -71C what will be on F?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
C/5 = (F – 32)/9
- 71 * 9/5 = F – 32
F – 32 = – 127.8
F = – 95.8
Hope it helps
In case of any difficulty, please feel free to ask again.
Regards
Arun